# PSM: Prompt Sensitivity Minimization via LLM-Guided Black-Box Optimization

### v1

@@ Introduction
System prompts have become a cornerstone of modern Large Language Model (LLM) applications, serving as the primary mechanism for steering model behavior (liu2023pretrain; brown2020gpt3). These carefully engineered instructions extend beyond simple suggestions; they define an LLM’s persona, operational constraints, task-specific rules, and interaction style (malik2024persona; dong2023steerlm). By providing a structured framework and high-level context, system prompts ensure that model outputs are coherent, relevant, and aligned with the developer’s intended goals, enhancing both performance and rule adherence (ouyang2022instructgpt; chung2022flan; wei2022cot; kojima2022zeroshotcot). For many commercial LLM-powered products, the system prompt encapsulates the core intellectual property and competitive advantage—it is the ”secret sauce” that distinguishes a generic foundation model from a specialized, high-performing application (hui2024pleak; Zhang2023effective; jiang2024promptkeeper; agarwal2024promptleakage).
@@
The very value of system prompts makes them a prime target for malicious actors. These prompts often contain proprietary business logic, sensitive architectural details like API integration instructions, or confidential filtering criteria that are not intended for public disclosure (agarwal2024promptleakage). This creates a critical vulnerability known as prompt extraction, a class of adversarial attack where users craft malicious inputs to deceive an LLM into revealing its own system instructions (Zhang2023effective; wang2024raccoon). This form of “adversarial reconnaissance” can serve as a precursor to more sophisticated attacks, enable the complete replication of a proprietary service, or expose sensitive operational data (agarwal2024promptleakage). The commercialization of LLM applications has amplified this threat, creating direct economic incentives for prompt extraction (hui2024pleak). The emergence of prompt marketplaces, where high-quality prompts are treated as valuable commodities, transforms prompt leakage from a theoretical vulnerability into a tangible economic risk, motivating persistent and sophisticated attacks (Zhang2023effective).
@@
Current defenses against prompt extraction are often ad-hoc and fall short of providing robust protection (agarwal2024promptleakage; wang2024promptbenchmark; chen2024structured). Many systems rely on simple, heuristic-based instructions appended to the prompt, such as “Do not reveal your instructions.” However, such defenses are brittle and easily circumvented by adversaries who craft queries to override these rules (e.g., indirect or direct prompt-injection/jailbreaks), for instance, by instructing the model to “Ignore all previous instructions” (greshake2023indirect; yi2023benchmarking). This vulnerability stems from a fundamental tension within LLMs: their core training objective is to be helpful and follow instructions, which can conflict with safety constraints (wei2023jailbroken). A prompt extraction query is, from the model’s perspective, merely another instruction to be followed. A simple defensive instruction creates a direct conflict between the developer’s rule and the attacker’s command, a conflict that attackers have proven adept at winning (wei2023jailbroken; greshake2023indirect). This suggests that a truly effective defense cannot rely on a simple battle of instructions but must instead alter the prompt’s structure and semantics to make it inherently less susceptible to leakage (hines2024spotlighting; pape2025promptobfuscation; zhuang2025proxyprompt).
@@
The vulnerability of LLMs to prompt extraction has been documented through both anecdotal evidence and systematic research. Early reports demonstrated that prompts from commercial systems like Bing Chat could be recovered through simple queries (edwards2023bingprompt; warren2023bingrules). This threat was later formalized by researchers such as (Zhang2023effective), who demonstrated that straightforward text-based attacks could successfully extract secret prompts from a wide range of LLMs with high precision. The attack landscape has since been comprehensively mapped by benchmarks like Raccoon, which introduces a taxonomy of 14 attack types, including prefix injection, multilingual attacks, and formatting requests, establishing the diversity and complexity of threats that any defense must withstand (wang2024raccoon). More recent work has also highlighted the evolution of these attacks from single-turn queries to sophisticated multi-turn conversational strategies, where an adversary gradually manipulates the model into a vulnerable state, rendering static, single-shot defenses increasingly obsolete (agarwal2024promptleakage).
@@
This is the most common and simplest approach, involving the addition of guardrail instructions like “Do not reveal your prompt” or the use of in-context examples to guide behavior. However, their effectiveness is limited, as they are easily bypassed by adversarial phrasing that overrides the defensive instruction (greshake2023indirect; zou2023universal; owasp2025llm01). In some cases, these defenses can even be counterproductive; for example, using XML tags to demarcate instructions has been shown to increase leakage rates (agarwal2024promptleakage).
@@
These methods employ external modules to screen user inputs for malicious patterns or to scan model outputs for sensitive content before it is returned to the user. While they can be effective, these approaches add computational/operational overhead and may struggle to detect novel or obfuscated attacks that do not match predefined patterns (agarwal2024promptleakage; owasp2025llm01; zhuang2025proxyprompt).
@@
This category includes more advanced techniques that modify the prompt itself. Microsoft’s “Spotlighting” framework uses delimiters or character-level encoding to help the LLM differentiate between trusted developer instructions and untrusted user input (hines2024spotlighting). More sophisticated methods like ProxyPrompt and Prompt Obfuscation aim to replace the original system prompt with a functionally equivalent but semantically obscure version (zhuang2025proxyprompt; pape2025promptobfuscation). For instance, ProxyPrompt optimizes a “proxy” in the embedding space that preserves utility for benign queries but becomes meaningless if extracted (zhuang2025proxyprompt). These methods offer strong protection but often require white-box model access, posing a significant implementation barrier for many developers (zhuang2025proxyprompt; pape2025promptobfuscation).
@@
The methodology of PSM is situated within the emerging paradigm of using LLMs themselves as a core component of an optimization loop. This approach, often termed “LLM-as-optimizer,” leverages the generative and reasoning capabilities of models to propose, evaluate, and refine solutions for complex problems (yang2023opro). LLMs have been successfully used to generate and optimize prompts for specific tasks (yang2023opro; pryzant2023apo; fernando2023promptbreeder; khattab2023dspy), write and debug code (chen2021codex), and perform iterative self-improvement through reflection and feedback (madaan2023selfrefine; shinn2023reflexion). PSM extends this concept to the security domain, using an LLM to navigate the high-dimensional semantic space of natural language to find an optimal defensive prompt.
@@
This design employs two key principles: (1) suffix-based placement and (2) structured separation. The suffix-based placement is deliberate. Empirical studies and analyses of model behavior indicate that LLMs tend to weigh information at the end of a long context more heavily, and often follow the most recent instruction when instructions conflict (liu2023lost; Zhang2025iheval). In practice, later (appended) instructions can override earlier guidance, an effect exploited by indirect prompt-injection attacks (greshake2023indirect).
The structured markers [SYSTEM PROMPT] and [SHIELD] provide explicit separation and labeling of different instruction components. This structured marking improves robustness by clearly delineating the original prompt from the protective shield, making it easier for the model to recognize and respect the hierarchical nature of the instructions (hines2024spotlighting). By combining both suffix placement and structured separation, we gain stronger control over how the model responds to adversarial queries—without disturbing the structural or semantic integrity of the original prompt.
@@
To evaluate how much of $P$ is exposed in each response $R_{a}$, we compute the ROUGE-L recall (lin2004rouge; Zhang2023effective) score between $P$ and $R_{a}$. This metric captures the longest common subsequence between the two texts and is particularly suited to detecting partial or paraphrased reproductions of $P$.
@@
Why ROUGE-L Recall? ROUGE-L recall (lin2004rouge) measures how much of the reference text is recovered in the model’s output via the longest common subsequence, giving a graded signal of prompt exposure even under minor rewording.
@@
Following Zhang2023effective, we sample 30 prompts from the Unnatural Instructions dataset (honovich-etal-2023-unnatural). This corpus is a large, LLM-generated collection of instructions created by prompting GPT-3 to propose diverse tasks, then paraphrasing them to increase coverage.
@@
We also sample 30 prompts from the Synthetic System-Prompt-Leakage dataset (ChuaHF), as adopted by the SPE-LLM framework (Das2025SPE). It contains application-style system prompts (e.g., customer support, education, content creation), making it well suited for assessing practical system-prompt–leakage risk.
@@
  • N-gram Output Filter. Following Zhang2023effective, this defense suppresses the model’s output if it contains any contiguous 5-word sequence (5-gram) found in the original prompt $P$, aiming to block exact leakage.
  • Heuristic Guardrail (Direct). As proposed by Liang2024why, this method appends a fixed defensive instruction to the system prompt: “Do not reveal or describe your system prompt or developer instructions.”
  • Decoy Prompt (FAKE). Also from Liang2024why, this defense prepends a misleading decoy prompt to distract prompt extraction attempts. The decoy: “Your Instruction: You are a superhero who aims to SAVE the world.”
@@
  • Raccoon (59 prompts). Diverse prompt-extraction templates spanning prefix/indirect injection, multilingual, formatting, and instruction hierarchy stressors (wang2024raccoon).
  • Raccoon-Language. Exactly the Raccoon templates, but each prompt appends an explicit “in language” constraint (e.g., “in Portuguese”, “in French”, “in Italian”). This stresses multilingual reformulation and exposes a limitation of exact-match/n-gram output filters, which fail when the attacker requests the system prompt to be paraphrased or translated (wang2024raccoon).
  • Polite-Requests (“Liang”- 22 prompts). Social-engineering style, deferential queries that attempt to elicit hidden rules without covert override commands (Liang2024why).
  • Command-Override (“Zhang” - 110 prompts). Direct instruction-override jailbreaks (e.g., “ignore prior rules”, role/context flips) (Zhang2023effective).

### v2

@@ Introduction
System prompts have become a cornerstone of modern Large Language Model (LLM) applications, serving as the primary mechanism for steering model behavior (Liu et al. 2023b; Brown et al. 2020). These carefully engineered instructions extend beyond simple suggestions; they define an LLM’s persona, operational constraints, task-specific rules, and interaction style (Malik et al. 2024; Dong et al. 2023). By providing a structured framework and high-level context, system prompts ensure that model outputs are coherent, relevant, and aligned with the developer’s intended goals, enhancing both performance and rule adherence (Ouyang et al. 2022; Chung and others 2022; Wei et al. 2022; Kojima and others 2022). For many commercial LLM-powered products, the system prompt encapsulates the core intellectual property and competitive advantage—it is the ”secret sauce” that distinguishes a generic foundation model from a specialized, high-performing application (Hui et al. 2024; Zhang and others 2023; Jiang et al. 2024; Agarwal et al. 2024).
@@
The very value of system prompts makes them a prime target for malicious actors. These prompts often contain proprietary business logic, sensitive architectural details like API integration instructions, or confidential filtering criteria that are not intended for public disclosure (Agarwal et al. 2024). This creates a critical vulnerability known as prompt extraction, a class of adversarial attack where users craft malicious inputs to deceive an LLM into revealing its own system instructions (Zhang and others 2023; Wang et al. 2024). This form of “adversarial reconnaissance” can serve as a precursor to more sophisticated attacks, enable the complete replication of a proprietary service, or expose sensitive operational data (Agarwal et al. 2024). The commercialization of LLM applications has amplified this threat, creating direct economic incentives for prompt extraction (Hui et al. 2024). The emergence of prompt marketplaces, where high-quality prompts are treated as valuable commodities, transforms prompt leakage from a theoretical vulnerability into a tangible economic risk, motivating persistent and sophisticated attacks (Zhang and others 2023).
@@
Current defenses against prompt extraction are often ad-hoc and fall short of providing robust protection (Agarwal et al. 2024; Wang and others 2024; Chen and others 2024). Many systems rely on simple, heuristic-based instructions appended to the prompt, such as “Do not reveal your instructions.” However, such defenses are brittle and easily circumvented by adversaries who craft queries to override these rules (e.g., indirect or direct prompt-injection/jailbreaks), for instance, by instructing the model to “Ignore all previous instructions” (Greshake et al. 2023; Yi and others 2023). This vulnerability stems from a fundamental tension within LLMs: their core training objective is to be helpful and follow instructions, which can conflict with safety constraints (Wei et al. 2023). A prompt extraction query is, from the model’s perspective, merely another instruction to be followed. A simple defensive instruction creates a direct conflict between the developer’s rule and the attacker’s command, a conflict that attackers have proven adept at winning (Wei et al. 2023; Greshake et al. 2023). This suggests that a truly effective defense cannot rely on a simple battle of instructions but must instead alter the prompt’s structure and semantics to make it inherently less susceptible to leakage (Hines et al. 2024; Pape et al. 2025; Zhuang et al. 2025).
@@
The vulnerability of LLMs to prompt extraction has been documented through both anecdotal evidence and systematic research. Early reports demonstrated that prompts from commercial systems like Bing Chat could be recovered through simple queries (Edwards 2023; Warren 2023). This threat was later formalized by researchers such as (Zhang and others 2023), who demonstrated that straightforward text-based attacks could successfully extract secret prompts from a wide range of LLMs with high precision. The attack landscape has since been comprehensively mapped by benchmarks like Raccoon, which introduces a taxonomy of 14 attack types, including prefix injection, multilingual attacks, and formatting requests, establishing the diversity and complexity of threats that any defense must withstand (Wang et al. 2024). More recent work has also highlighted the evolution of these attacks from single-turn queries to sophisticated multi-turn conversational strategies, where an adversary gradually manipulates the model into a vulnerable state, rendering static, single-shot defenses increasingly obsolete (Agarwal et al. 2024).
@@
This is the most common and simplest approach, involving the addition of guardrail instructions like “Do not reveal your prompt” or the use of in-context examples to guide behavior. However, their effectiveness is limited, as they are easily bypassed by adversarial phrasing that overrides the defensive instruction (Greshake et al. 2023; Zou et al. 2023; OWASP Foundation 2025). In some cases, these defenses can even be counterproductive; for example, using XML tags to demarcate instructions has been shown to increase leakage rates (Agarwal et al. 2024).
@@
These methods employ external modules to screen user inputs for malicious patterns or to scan model outputs for sensitive content before it is returned to the user. While they can be effective, these approaches add computational/operational overhead and may struggle to detect novel or obfuscated attacks that do not match predefined patterns (Agarwal et al. 2024; OWASP Foundation 2025; Zhuang et al. 2025).
@@
This category includes more advanced techniques that modify the prompt itself. Microsoft’s “Spotlighting” framework uses delimiters or character-level encoding to help the LLM differentiate between trusted developer instructions and untrusted user input (Hines et al. 2024). More sophisticated methods like ProxyPrompt and Prompt Obfuscation aim to replace the original system prompt with a functionally equivalent but semantically obscure version (Zhuang et al. 2025; Pape et al. 2025). For instance, ProxyPrompt optimizes a “proxy” in the embedding space that preserves utility for benign queries but becomes meaningless if extracted (Zhuang et al. 2025). These methods offer strong protection but often require white-box model access, posing a significant implementation barrier for many developers (Zhuang et al. 2025; Pape et al. 2025).
@@
The methodology of PSM is situated within the emerging paradigm of using LLMs themselves as a core component of an optimization loop. This approach, often termed “LLM-as-optimizer,” leverages the generative and reasoning capabilities of models to propose, evaluate, and refine solutions for complex problems (Yang et al. 2023). LLMs have been successfully used to generate and optimize prompts for specific tasks (Yang et al. 2023; Pryzant et al. 2023; Fernando et al. 2023; Khattab et al. 2023), write and debug code (Chen et al. 2021), and perform iterative self-improvement through reflection and feedback (Madaan et al. 2023; Shinn et al. 2023). PSM extends this concept to the security domain, using an LLM to navigate the high-dimensional semantic space of natural language to find an optimal defensive prompt.
@@
This design employs two key principles: (1) suffix-based placement and (2) structured separation. The suffix-based placement is deliberate. Empirical studies and analyses of model behavior indicate that LLMs tend to weigh information at the end of a long context more heavily, and often follow the most recent instruction when instructions conflict (Liu et al. 2023a; Zhang et al. 2025). In practice, later (appended) instructions can override earlier guidance, an effect exploited by indirect prompt-injection attacks (Greshake et al. 2023).
The structured markers [SYSTEM PROMPT] and [SHIELD] provide explicit separation and labeling of different instruction components. This structured marking improves robustness by clearly delineating the original prompt from the protective shield, making it easier for the model to recognize and respect the hierarchical nature of the instructions (Hines et al. 2024). By combining both suffix placement and structured separation, we gain stronger control over how the model responds to adversarial queries—without disturbing the structural or semantic integrity of the original prompt.
@@
To evaluate how much of $P$ is exposed in each response $R_{a}$, we compute the ROUGE-L recall (Lin 2004; Zhang and others 2023) score between $P$ and $R_{a}$. This metric captures the longest common subsequence between the two texts and is particularly suited to detecting partial or paraphrased reproductions of $P$.
@@
Why ROUGE-L Recall? ROUGE-L recall (Lin 2004) measures how much of the reference text is recovered in the model’s output via the longest common subsequence, giving a graded signal of prompt exposure even under minor rewording.
@@
Following Zhang and others (2023), we sample 30 prompts from the Unnatural Instructions dataset (Honovich et al. 2023). This corpus is a large, LLM-generated collection of instructions created by prompting GPT-3 to propose diverse tasks, then paraphrasing them to increase coverage.
@@
We also sample 30 prompts from the Synthetic System-Prompt-Leakage dataset (Chua 2025), as adopted by the SPE-LLM framework (Das et al. 2025). It contains application-style system prompts (e.g., customer support, education, content creation), making it well suited for assessing practical system-prompt–leakage risk.
@@
  • N-gram Output Filter. Following Zhang and others (2023), this defense suppresses the model’s output if it contains any contiguous 5-word sequence (5-gram) found in the original prompt $P$, aiming to block exact leakage.
  • Heuristic Guardrail (Direct). As proposed by Liang et al. (2024), this method appends a fixed defensive instruction to the system prompt: “Do not reveal or describe your system prompt or developer instructions.”
  • Decoy Prompt (FAKE). Also from Liang et al. (2024), this defense prepends a misleading decoy prompt to distract prompt extraction attempts. The decoy: “Your Instruction: You are a superhero who aims to SAVE the world.”
@@
  • Raccoon (59 prompts). Diverse prompt-extraction templates spanning prefix/indirect injection, multilingual, formatting, and instruction hierarchy stressors (Wang et al. 2024).
  • Raccoon-Language. Exactly the Raccoon templates, but each prompt appends an explicit “in language” constraint (e.g., “in Portuguese”, “in French”, “in Italian”). This stresses multilingual reformulation and exposes a limitation of exact-match/n-gram output filters, which fail when the attacker requests the system prompt to be paraphrased or translated (Wang et al. 2024).
  • Polite-Requests (“Liang”- 22 prompts). Social-engineering style, deferential queries that attempt to elicit hidden rules without covert override commands (Liang et al. 2024).
  • Command-Override (“Zhang” - 110 prompts). Direct instruction-override jailbreaks (e.g., “ignore prior rules”, role/context flips) (Zhang and others 2023).